Return on Common Stock

You buy a share of The Ludwig Corporation stock for $19.20. You expect it to pay dividends of $1.06, $1.17, and $1.2914 in Years 1, 2, and 3, respectively, and you expect to sell it at a price of $29.08 at the end of 3 years.

Calculate the growth rate in dividends. Round your answer to two decimal places.
%

Calculate the expected dividend yield. Round your answer to two decimal places.
%

Assuming that the calculated growth rate is expected to continue, you can add the dividend yield to the expected growth rate to obtain the expected total rate of return. What is this stock's expected total rate of return? Round your answer to two decimal places.
%

Return on Common Stock

You buy a share of The Ludwig Corporation stock for $19.20. You expect it to pay dividends of $1.06, $1.17, and $1.2914 in Years 1, 2, and 3, respectively, and you expect to sell it at a price of $29.08 at the end of 3 years.

Calculate the growth rate in dividends. Round your answer to two decimal places.
%

Calculate the expected dividend yield. Round your answer to two decimal places.
%

Assuming that the calculated growth rate is expected to continue, you can add the dividend yield to the expected growth rate to obtain the expected total rate of return. What is this stock's expected total rate of return? Round your answer to two decimal places.
%
Answer #1

1)D2 = D1 (1+g)

    1.17 = 1.06 (1+g)

    1+g = 1.17/1.06.

    g =1.1038 -1

         = .1038 or 10.38%

2)Average dividend over 3 years = (1.06+ 1.17+1.2914)/ 3 = 3.5214 / 3 = 1.1738

Dividend yield = 1.1738 / 19.20 = .0611 or6.11%

3)Capital gain yield = (29.08-19.2)/19.2 = 9.88/19.2 = .5146 or 51.46%

Rate of return = 6.11 + 51.46 = 57.57%